# The Stepanov theorem for Q-valued functions

### v1

@@ 2.2. Differentiaition and completeness
Let $A$ be a nonempty subset of $X$, $\mathcal{D}$ be an $X$-distribution and let ${f}:{A}\rightarrow{M}$ be an arbitrary function. For every $x\in A$ accumulation point we say that $f$ is $\mathcal{D}$-differentiable at $x$ if and only if there exists $[{F}:{B}\rightarrow{M}]\in\mathcal{D}_{x,f(x)}$ such that
$\lim_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[f(y),F(y)\right]}{\mathtt{J}\left(y,x\right)}=0\text{.}$ (2.4)
@@
$\mu-\lim_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[f(y),F(y)\right]}{\mathtt{J}\left(y,x\right)}=0$
@@
$\left\lvert\nabla f\right\rvert^{+}(x)=\limsup_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[f(y),f(x)\right]}{\mathtt{J}\left(y,x\right)}\\ \leq\limsup_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[f(y),F(y)\right]}{\mathtt{J}\left(y,x\right)}+\limsup_{y\to x}\frac{d\left[F(y),F(x)\right]}{\mathtt{J}\left(y,x\right)}=\left\lvert F\right\rvert_{x,f(x)}\text{.}$
Next we prove that if $x$ is a $\mu$-interior point for $A$ then $f$ has exactly one $\mathcal{D}$-differential at $x$ and equality holds in the last inequality. Let $[G]\in\mathcal{D}_{x,f(x)}$ be another $\mathcal{D}$-differential of $f$ at $x$, then we immediately get
$\lim_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}=0$
@@
$\mu-\limsup_{y\to x}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}=\mu-\lim_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}\\ \leq\lim_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}=0\text{.}$
@@
$\left\lvert T_{x}f\right\rvert_{x,f(x)}=\mu-\limsup_{y\to x}\frac{d[F(y),F(x)]}{\mathtt{J}(y,x)}\\ =\mu-\limsup_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d[F(y),F(x)]}{\mathtt{J}(y,x)}\leq\left\lvert\nabla f\right\rvert^{+}(x)\text{.}\qed$
In a similar fashion we can prove the following:
@@
We then say that a function ${f}:{A}\rightarrow{M}$ defined on a nonempty set $A\subseteq X$ is a.e. $\mathcal{D}$-differentiable on $A_{f}$ if and only if on every $\mu$-measurable subset $B\subseteq A_{f}$ $\mu$-measurable we have that $f$ is $\mathcal{D}$-differentiable for a.e. point of $B$. In the same way we define functions a.e. approximately $\mathcal{D}$-differentiable on $A_{f}^{\mu}$.
@@
for every closed nonempty subset $C\subseteq X$ (with respect to the $\mathtt{J}$-topology) and every Lipschitz function ${g}:{C}\rightarrow{M}$ we have that $g$ is a.e. approximately $\mathcal{D}$-differentiable on $A_{f}^{\mu}$.
@@
For Lipschitz functions $\mathcal{D}$-differentiability and approximate $\mathcal{D}$-differentiability are equivalent for a.e. point of $C$, where $\mathcal{D}$ is an $X$-distribution. Moreover, both $A_{f}$ and $A_{f}^{\mu}$ are Borel sets with $\mu\left({A_{f}}\Delta{A_{f}^{\mu}}\right)$ $=0$.
@@
    if $\mathcal{D}$ is differentiably complete then every $\mathcal{D}\subseteq\mathcal{C}$ is also differentiably complete;
@@
Of these statements, only point 3 is not trivial, however its proof is quite simple thanks to uniqueness results contained in propositions 2.6 and 2.7. Let then ${f}:{C}\rightarrow{M}$ be a Lipschitz function defined on a closed subset $C$ of $X$, then there exists a measurable subset $V={V}^{1}$ of $C$ such that for every $i\in{\mathbb{N}}$ $f$ is approximately $\mathcal{D}^{i}$-differentiable at $x$ with differential $\left[F_{i}\right]\in\mathcal{D}^{i}$. By uniqueness we have $\left[F_{i}\right]=\left[F_{1}\right]=T_{x}f$ for every $i\in{\mathbb{N}}$, therefore $T_{x}f\in\bigcap^{\infty}_{i=1}\mathcal{D}^{i}_{x,f(x)}=\mathcal{D}_{x,f(x)}$.
@@
A well-know result proven for the first time by Cheeger ([2]) states that under some conditions on the base space $X$ (in particular, the existence of a $p$-Poincaré type inequality for some $1\leq p<+\infty$) then $\mathcal{D}$ is a complete distribution for every $k\in{\mathbb{N}}$ and $\sup N_{\alpha}<+\infty$, in other words the metric space $X$ is “locally finite dimensional”.
In the last part of this section we will prove the equivalence of (RC) with a $\mathcal{D}$-version of the classical Stepanov theorem for differentiability and approximate differentiability. We first need to generalize proposition 2.3.
@@
    let $E\subseteq X$ be a nonempty set such that $x\in{\left(\overline{E}\right)}^{1}$ and for every $y\in E$ let $B_{y}$ be a measurable set so that for every $z\in\overline{E}\setminus\left\{{x}\right\}$ sufficiently close to $x$ there exist $l,\eta>0$ such that $B_{y}$ is $(\delta,\eta)$-full at $y$ for every $y\in B_{l}(z)$, then there exist a $\mu$-neighborhood $U$ of $x$ and an application ${r_{x}}:{U}\rightarrow{E}$ such that $y\in B_{r_{x}(y)}$ for every $y\in U\setminus\left\{{x}\right\}$.
@@
Fix $\delta>0$ and assume by contradiction that 1. does not hold. Thus $\mu\left[B_{y}\cap B_{t}(y)\right]\leq\mu\left[B_{3t}(y)\setminus B_{y}\right]$ with $3t<\eta$ and then
$\begin{split}(1-\delta)\mu\left[B_{t}(x)\right]&\leq\mu\left[B_{t}(x)\cap B_{x}\right]\\ &\leq\mu\left[B_{3t}(y)\setminus B_{y}\right]\\ &\leq\delta\mu\left[B_{3t}(y)\right]\\ &\leq\delta\mathcal{L}(K,4)\mu\left[B_{t}(x)\right]\end{split}$
@@
To prove 2 we can assume without loss of generality that $x$ is a $\mu$-accumulation point, that is $\mu(\left\{{x}\right\})=0$. Notice also that we can find a sequence $y_{n}\in E$ such that $\overline{E}=\overline{\left\{{y_{n}}\mathrel{}\middle|\mathrel{}{n\in{\mathbb{N}}}\right\}}$. Then we need only to prove that
@@
Otherwise we can find $Z\subseteq\overline{E}$ sufficiently near $x$ such that $Z\cap B_{y_{n}}=\emptyset$ and
@@
for some $z\in Z\setminus\left\{{x}\right\}$ and $0<\eta_{0}$, in particular $Z$ if $(\delta,\eta_{0})$-full at $z$. By assumption we can find $l>0$ and $0<\eta<\eta_{0}$ such that $B_{y_{n}}$ is $(\delta,\eta)$-full at $y_{n}$ for every $y_{n}\in B_{l}(z)$. Moreover, for $n$ large we also have $\mathtt{J}\left(y_{n},z\right)<\eta/3$ therefore by 1. we would have $Z\cap B_{y_{n}}\neq\emptyset$, a contradiction. ∎
@@
Conversely let $\mathcal{D}$ be an approximate $X$-distribution such that every function ${f}:{X}\rightarrow{M}$ is approximately $\mathcal{D}$-differentiable on $A_{f}^{\mu}$, then $\mathcal{D}$ is differentiably complete.
We first prove that if $\mathcal{D}$ is complete then ${f}:{A}\rightarrow{M}$ is a.e. approximately $\mathcal{D}$-differentiable on every measurable subset $B$ of $\mathcal{A}_{f}^{\mu}$. Fix a point $P\in M$ and for every $i,j\in{\mathbb{N}}$ let $E_{ij}\subseteq A$ such that $x\in E_{ij}$ if and only if
@@
Since $M$ is complete we can uniquely extend $f_{ij}$ on $\overline{E_{ij}}$, therefore there exists a measurable subset $\tilde{G}_{ij}$ of $\overline{E_{ij}}$ such that $\mu\left(\overline{E_{ij}}\setminus\tilde{G}_{ij}\right)=0$ and $f_{ij}$ is approximately $\mathcal{D}$-differentiable on each point of $\tilde{G}_{ij}$.
@@
and $f_{ij}$ is approximately $\mathcal{D}$-differentiable on every point of $G_{ij}$. Let then $x\in G_{ij}$ and $\left[F_{ij}\right]\in\mathcal{D}_{x,f(x)}$ be the differential of $f_{ij}$ in $x$, then for every $\varepsilon>0$ there exists a measurable $\mu$-neighborhood $U_{x}\subseteq X$ of $x$ such that $d\left[f_{ij}(y),F_{ij}(y)\right]<\varepsilon\mathtt{J}(y,z)$ for every $y\in U_{x}\cap\overline{E_{ij}}$.
Let ${s}:{U}\rightarrow{E_{ij}}$ be any function defined on a $\mu$-neighborhood of $x$ such that $\mathtt{J}\left[s(y),x\right]\leq L\mathtt{J}(y,x)$ near $x$. Therefore,
$\begin{split}d\left[f(y),F_{ij}(y)\right]&\leq d\left[f(y),f\left[s_{x}(y)\right]\right]+d\left[f_{ij}\left[s_{x}(y)\right],F_{ij}\left[s_{x}(y)\right]\right]\\ &\qquad+d\left[F_{ij}\left(y\right),F_{ij}\left[s_{x}(y)\right]\right]\\ &\leq d\left[f(y),f\left[s_{x}(y)\right]\right]+o\left(\mathtt{J}\left[s_{x}(y),x\right]\right)+K\mathtt{J}\left(s_{x}(y),y\right)\\ &=d\left[f(y),f\left[s_{x}(y)\right]\right]+o\left(\mathtt{J}\left(y,x\right)\right)\text{.}\end{split}$
@@
Clearly each $\tilde{B}_{y}$ is $\left[\delta,R(y)\mathtt{J}(y,x)\right]$-full at $y$ and $R(y)$ depends only on $\mathtt{J}(y,x)$, therefore assumptions at point 2. in lemma 2.10 are satisfied and there exists ${s}:{U}\rightarrow{E_{ij}}$ such that $y\in\tilde{B}_{s(y)}$ for $y$ near $x$. In particular, $y\in B_{s(y)}$ then
$d[f(y),f[s(y)]]\leq i\mathtt{J}\left[y,s(y)\right]\text{.}$
Moreover, since $R(y)<1/2$, we also have $\mathtt{J}\left[y,s(y)\right]=o\left(\mathtt{J}(y,x)\right)$ therefore
$d[f(y),f[s(y)]]=o(\mathtt{J}(y,x))$
and $f$ is approximately $\mathcal{D}$-differentiable at $x$.
The converse if trivial. Indeed if every function ${f}:{X}\rightarrow{M}$ is approximately $\mathcal{D}$-differentiable a.e. on $A_{f}^{\mu}$ then to prove that $\mathcal{D}$ is complete we need only to accordingly extend any Lipschitz function ${g}:{C}\rightarrow{M}$ on all $X$ with $C$ closed. If $\tilde{g}$ is any extension of $g$ on all $X$ then we always have
@@
therefore $g$ is automatically a.e. approximately $\mathcal{D}$-differentiable, then $\mathcal{D}$ is complete.
@@
Conversely let $\mathcal{D}$ be an $X$-distribution such that every function ${f}:{X}\rightarrow{M}$ is $\mathcal{D}$-differentiable on $A_{f}$, then $\mathcal{D}$ is differentiably complete.
@@
Now assume that every function ${f}:{X}\rightarrow{M}$ is $\mathcal{D}$-differentiable a.e. on $A_{f}$, we will prove that $\mathcal{D}$ is complete or equivalently that every Lipschitz function ${g}:{C}\rightarrow{M}$ defined on any closed subset $C\subseteq X$ is $\mathcal{D}$-differentiable a.e. on $C$. We want to define an extension of $g$ to whole $X$, an even though we cannot expect that such extension is Lipschitz it has useful properties:
Proposition 2.13.
If ${f}:{C}\rightarrow{M}$ is a generic function defined on a nonempty closed subset $C$ of $X$ then there exists ${\tilde{f}}:{X}\rightarrow{M}$ such that $\left.{\tilde{f}}\right|_{C}=f$ and $A_{f}\subseteq A_{\tilde{f}}$.
Proof.
We define $\tilde{f}$ in the following way: if $x\in C$ then we set $\tilde{f}(x)=f(x)$, instead if $x\notin C$ then we set $\tilde{f}(x)=f(y)$ where $y\in C$ is any point that satisfies
@@
Clearly $\tilde{f}$ is an extension of $f$, we need only to prove that $A_{f}\subseteq A_{\tilde{f}}$. Let $x\in A_{f}\cap\partial C$, therefore there exist $K,\delta>0$ such that if $y\in C\cap B_{\delta}(x)$ then $d[f(x),f(y)]\leq K\mathtt{J}\left(x,y\right)$.
Let $z\in B_{\delta/2}(x)\setminus C$, we can then find $y\in C\cap B_{\delta}(x)$ such that $\mathtt{J}\left(y,z\right)<2\operatorname{dist}(z,C)$ and $\tilde{f}(z)=f(y)$. Therefore,
$d[\tilde{f}(x),\tilde{f}(z)]=d[f(x),f(y)]\leq K\mathtt{J}\left(x,y\right)\leq K\left(\mathtt{J}\left(x,z\right)+2\operatorname{dist}(z,C)\right)\\ \leq 3K\mathtt{J}\left(x,z\right)$
which implies $x\in A_{\tilde{f}}$ with $\left\lvert\nabla\tilde{f}\right\rvert^{+}(x)\leq 3\left\lvert\nabla f\right\rvert^{+}(x)$. ∎
Now let ${f}:{C}\rightarrow{M}$ be Lipschitz and $C\subseteq X$ closed, then we have an extension ${\tilde{f}}:{X}\rightarrow{M}$ such that $C\subseteq A_{\tilde{f}}$. Since $\tilde{f}$ is $\mathcal{D}$-differentiable a.e. on $A_{\tilde{f}}$ we have also that $g$ is $\mathcal{D}$-differentiable a.e. on $C$, therefore $\mathcal{D}$ is complete.

### v3

@@ 2.2. Differentiation and completeness
Let $A$ be a nonempty subset of $X$, $\mathcal{D}$ be an $X$-distribution, and let ${f}:{A}\rightarrow{M}$ be an arbitrary function. For every $x\in A$ accumulation point we say that $f$ is $\mathcal{D}$-differentiable at $x$ if and only if there exists $[{F}:{B}\rightarrow{M}]\in\mathcal{D}_{x,f(x)}$ such that
$\lim_{A\ni y\to x}\frac{d\left[f(y),F(y)\right]}{\mathtt{J}\left(y,x\right)}=0\text{.}$ (2.4)
@@
$\mu-\lim_{A\ni y\to x}\frac{d\left[f(y),F(y)\right]}{\mathtt{J}\left(y,x\right)}=0$
@@
$\left\lvert\nabla f\right\rvert^{+}(x)=\limsup_{A\ni y\to x}\frac{d\left[f(y),f(x)\right]}{\mathtt{J}\left(y,x\right)}\\ \leq\limsup_{A\ni y\to x}\frac{d\left[f(y),F(y)\right]}{\mathtt{J}\left(y,x\right)}+\limsup_{y\to x}\frac{d\left[F(y),F(x)\right]}{\mathtt{J}\left(y,x\right)}=\left\lvert F\right\rvert_{x,f(x)}\text{.}$
Next, we prove that if $x$ is a $\mu$-interior point for $A$ then $f$ has exactly one $\mathcal{D}$-differential at $x$ and equality holds in the last inequality. Let $[G]\in\mathcal{D}_{x,f(x)}$ be another $\mathcal{D}$-differential of $f$ at $x$, then we immediately get
$\lim_{A\ni y\to x}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}=0$
@@
$\mu-\limsup_{y\to x}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}=\mu-\lim_{A\ni y\to x}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}\\ \leq\lim_{A\ni y\to x}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}=0\text{.}$
@@
$\left\lvert T_{x}f\right\rvert_{x,f(x)}=\mu-\limsup_{y\to x}\frac{d[F(y),F(x)]}{\mathtt{J}(y,x)}\\ =\mu-\limsup_{A\ni y\to x}\frac{d[F(y),F(x)]}{\mathtt{J}(y,x)}\leq\left\lvert\nabla f\right\rvert^{+}(x)\text{.}\qed$
Similarly, we can prove the following:
@@
We then say that a function ${f}:{A}\rightarrow{M}$ defined on a nonempty set $A\subseteq X$ is a.e. $\mathcal{D}$-differentiable on $A_{f}$ if and only if on every $\mu$-measurable subset $B$ of $A_{f}$ we have that $f$ is $\mathcal{D}$-differentiable for a.e. point of $B$. In the same way, we define functions a.e. approximately $\mathcal{D}$-differentiable on $A_{f}^{\mu}$.
@@
for every closed nonempty subset $C\subseteq X$ (with respect to the $\mathtt{J}$-topology) and every Lipschitz function ${g}:{C}\rightarrow{M}$ we have that $g$ is a.e. approximately $\mathcal{D}$-differentiable on $A_{g}^{\mu}$.
@@
For Lipschitz functions defined on closed subsets $\mathcal{D}$-differentiability and approximate $\mathcal{D}$-differentiability are equivalent for a.e. point of $C$, where $\mathcal{D}$ is an $X$-distribution. Moreover, both $A_{g}$ and $A_{g}^{\mu}$ are Borel sets with $\mu\left({A_{g}}\Delta{A_{g}^{\mu}}\right)$ $=0$.
@@
    if $\mathcal{D}$ is differentiably complete then every $\mathcal{C}\supseteq\mathcal{D}$ is also differentiably complete;
@@
Of these statements, only (3) is not trivial, however its proof is quite simple thanks to uniqueness results contained in propositions 2.6 and 2.7. Let then ${f}:{C}\rightarrow{M}$ be a Lipschitz function defined on a closed subset $C$ of $X$, then there exists a measurable subset $V={V}^{1}$ of $C$ such that for every $i\in{\mathbb{N}}$ $f$ is approximately $\mathcal{D}^{i}$-differentiable at $x$ with differential $\left[F_{i}\right]\in\mathcal{D}^{i}$. By uniqueness we have $\left[F_{i}\right]=\left[F_{1}\right]=T_{x}f$ for every $i\in{\mathbb{N}}$, therefore $T_{x}f\in\bigcap^{\infty}_{i=1}\mathcal{D}^{i}_{x,f(x)}=\mathcal{D}_{x,f(x)}$.
@@
A well-known result proven for the first time by Cheeger ([Cheeger1999]) states that under some conditions on the base space $X$ (in particular, the existence of a $p$-Poincaré type inequality for some $1\leq p<+\infty$) then $\mathcal{D}$ is a complete distribution for every $k\in{\mathbb{N}}$ and $\sup N_{\alpha}<+\infty$, in other words the metric space $X$ is “locally finite dimensional”.
In the last part of this section, we will prove the equivalence of (RC) with a $\mathcal{D}$-version of the classical Stepanov theorem for differentiability and approximate differentiability. We first need to generalize proposition 2.3.
@@
    let $E\subseteq X$ be a nonempty set such that $x\in{\left(\overline{E}\right)}^{1}$ and for every $y\in E$ let $B_{y}$ be a measurable set so that for every $z\in\overline{E}\setminus\left\{{x}\right\}$ sufficiently close to $x$ there exist $s,\eta>0$ such that $B_{y}$ is $(\delta,\eta)$-full at $y$ for every $y\in B_{s}(z)$, then there exist a $\mu$-neighborhood $U$ of $x$ and a map ${r_{x}}:{U}\rightarrow{E}$ such that $y\in B_{r_{x}(y)}$ for every $y\in U\setminus\left\{{x}\right\}$.
@@
Fix $\delta>0$ and assume by contradiction that (1) does not hold. Thus $\mu\left(B_{x}\cap B_{t}(x)\right)\leq\mu\left(B_{3t}(y)\setminus B_{y}\right)$ with $3t<\eta$ and then
$\begin{split}(1-\delta)\mu\left(B_{t}(x)\right)&\leq\mu\left(B_{t}(x)\cap B_{x}\right)\\ &\leq\mu\left(B_{3t}(y)\setminus B_{y}\right)\\ &\leq\delta\mu\left(B_{3t}(y)\right)\\ &\leq\delta\mathcal{L}(K,4)\mu\left(B_{t}(x)\right)\end{split}$
@@
To prove (2) we can assume without loss of generality that $x$ is a $\mu$-accumulation point, that is $\mu(\left\{{x}\right\})=0$. Notice also that we can find a sequence $y_{n}\in E$ such that $\overline{E}=\overline{\left\{{y_{n}}\mathrel{}\middle|\mathrel{}{n\in{\mathbb{N}}}\right\}}$. Then we need only to prove that
@@
Otherwise we can find $Z\subseteq\overline{E}$ sufficiently near to $x$ such that $Z\cap B_{y_{n}}=\emptyset$ and
@@
for some $z\in Z\setminus\left\{{x}\right\}$ and $0<\eta_{0}$, in particular $Z$ is $(\delta,\eta_{0})$-full at $z$. By assumption we can find $s>0$ and $0<\eta<\eta_{0}$ such that $B_{y_{n}}$ is $(\delta,\eta)$-full at $y_{n}$ for every $y_{n}\in B_{s}(z)$. Moreover, for $n$ large we also have $\mathtt{J}\left(y_{n},z\right)<\eta/3$ therefore by (1) we would have $Z\cap B_{y_{n}}\neq\emptyset$, a contradiction. ∎
@@
Conversely let $\mathcal{D}$ be an approximate $X$-distribution such that every function ${f}:{X}\rightarrow{M}$ is a.e. approximately $\mathcal{D}$-differentiable on $A_{f}^{\mu}$, then $\mathcal{D}$ is differentiably complete.
Proof.
We first prove that if $\mathcal{D}$ is complete then ${f}:{A}\rightarrow{M}$ is a.e. approximately $\mathcal{D}$-differentiable on every measurable subset $B$ of $A_{f}^{\mu}$. Fix a point $P\in M$ and for every $i,j\in{\mathbb{N}}$ let $E_{ij}\subseteq A^{\mu}_{f}$ be such that $x\in E_{ij}$ if and only if
@@
Since $M$ is complete we can uniquely extend $f_{ij}$ to a Lipschitz function on $\overline{E_{ij}}$, and since $\mathcal{D}$ is complete there exists a measurable subset $\tilde{G}_{ij}$ of $A_{f_{ij}}^{\mu}\subseteq\overline{E_{ij}}$ such that $\mu\left(A_{f_{ij}}^{\mu}\setminus\tilde{G}_{ij}\right)=0$ and $f_{ij}$ is approximately $\mathcal{D}$-differentiable on each point of $\tilde{G}_{ij}$. Since $E_{ij}\subseteq A^{\mu}_{f}$, every point of $\overline{E_{ij}}$ is an accumulation point for $X$, thus $\mu\left(\overline{E_{ij}}\setminus A_{f_{ij}}^{\mu}\right)=0$.
@@
and $f_{ij}$ is approximately $\mathcal{D}$-differentiable on every point of $G_{ij}$. Let then $x\in G_{ij}$ and $\left[F_{ij}\right]\in\mathcal{D}_{x,f(x)}$ be the differential of $f_{ij}$ in $x$, then for every $\varepsilon>0$ there exists a measurable $\mu$-neighborhood $U_{x}\subseteq X$ of $x$ such that $d\left[f_{ij}(y),F_{ij}(y)\right]<\varepsilon\mathtt{J}(y,x)$ for every $y\in U_{x}\cap\overline{E_{ij}}$.
Let ${s}:{U}\rightarrow{E_{ij}}$ be any function defined on a $\mu$-neighborhood $U$ of $x$ such that $\mathtt{J}\left[s(y),x\right]\leq L\mathtt{J}(y,x)$ near $x$ and $\mathtt{J}\left[s(y),y\right]=o\left(\mathtt{J}(y,x)\right)$ for $\mathtt{J}(y,x)\to 0$. Therefore,
$\begin{split}d\left[f(y),F_{ij}(y)\right]&\leq d\left[f(y),f\left[s(y)\right]\right]+d\left[f_{ij}\left[s(y)\right],F_{ij}\left[s(y)\right]\right]\\ &\qquad+d\left[F_{ij}\left(y\right),F_{ij}\left[s(y)\right]\right]\\ &\leq d\left[f(y),f\left[s(y)\right]\right]+o\left(\mathtt{J}\left[s(y),x\right]\right)+K\mathtt{J}\left(s(y),y\right)\\ &=d\left[f(y),f\left[s(y)\right]\right]+o\left(\mathtt{J}\left(y,x\right)\right)\text{.}\end{split}$
@@
Clearly each $\tilde{B}_{y}$ is $\left[\delta,R(y)\mathtt{J}(y,x)\right]$-full at $y$ (since $\mathtt{J}(y,x)<1/j$) and $R(y)$ depends only on $\mathtt{J}(y,x)$, therefore assumptions at point (2) in lemma 2.10 are satisfied and there exists ${s}:{U}\rightarrow{E_{ij}}$ such that $y\in\tilde{B}_{s(y)}$ for $y$ near $x$.
In particular, $y\in B_{s(y)}$ and $\mathtt{J}(y,s(y))\leq R\left(s(y)\right)\mathtt{J}(y,x)<\frac{1}{2}\mathtt{J}(y,x)$ which leads to $\mathtt{J}(s(y),x)<\frac{3}{2}\mathtt{J}(y,x)$. Moreover, $\mathtt{J}(y,s(y))\leq\mathtt{J}(y,x)^{2}=o\left(\mathtt{J}(y,x)\right)$ when $\mathtt{J}(y,x)\to 0$, therefore
$d[f(y),f[s(y)]]\leq i\mathtt{J}\left[y,s(y)\right]=o(\mathtt{J}(y,x))$
because $s(y)\in E_{ij}$ and $y\in B_{s(y)}$, thus $f$ is approximately $\mathcal{D}$-differentiable at $x$.
The converse is trivial. Indeed if every function ${f}:{X}\rightarrow{M}$ is approximately $\mathcal{D}$-differentiable a.e. on $A_{f}^{\mu}$ then to prove that $\mathcal{D}$ is complete we need only to accordingly extend any Lipschitz function ${g}:{C}\rightarrow{M}$ on all $X$ with $C$ closed. If $\tilde{g}$ is any extension of $g$ on all $X$ then we always have
@@
therefore $g$ is automatically a.e. approximately $\mathcal{D}$-differentiable, then $\mathcal{D}$ is complete. ∎
@@
Conversely let $\mathcal{D}$ be an $X$-distribution such that every function ${f}:{X}\rightarrow{M}$ is a.e. $\mathcal{D}$-differentiable on $A_{f}$, then $\mathcal{D}$ is differentiably complete.
Proof.
@@
Now assume that every function ${f}:{X}\rightarrow{M}$ is a.e. $\mathcal{D}$-differentiable on $A_{f}$, we will prove that $\mathcal{D}$ is complete or equivalently that every Lipschitz function ${g}:{C}\rightarrow{M}$ defined on any closed subset $C\subseteq X$ is a.e. $\mathcal{D}$-differentiable on $C$. We will find an extension $\tilde{g}$ of $g$ to whole $X$, and even though we cannot expect that such extension is Lipschitz it has still some useful properties. More precisely, for every map ${g}:{C}\rightarrow{M}$ defined on a closed set $C\subseteq X$ we can always find an extension ${\tilde{g}}:{X}\rightarrow{M}$ such that $A_{g}\subseteq A_{\tilde{g}}$.
We define $\tilde{g}$ in the following way: if $x\in C$ then we set $\tilde{g}(x)=g(x)$, instead if $x\notin C$ then we set $\tilde{g}(x)=g(y)$ where $y\in C$ is any point that satisfies
@@
Let $x\in A_{g}\cap\partial C$, therefore there exist $K,\delta>0$ such that if $y\in C\cap B_{\delta}(x)$ then $d\left[\tilde{g}(x),\tilde{g}(y)\right]=d\left[g(x),g(y)\right]\leq K\mathtt{J}\left(x,y\right)$.
Let $z\in B_{\delta/3}(x)\setminus C$, we can then find $y\in C\cap B_{\delta}(x)$ such that $\mathtt{J}\left(y,z\right)<2\operatorname{dist}(z,C)$ and $\tilde{g}(z)=g(y)$. Therefore,
$d[\tilde{g}(x),\tilde{g}(z)]=d[g(x),g(y)]\leq K\mathtt{J}\left(x,y\right)\leq K\left(\mathtt{J}\left(x,z\right)+2\operatorname{dist}(z,C)\right)\\ \leq 3K\mathtt{J}\left(x,z\right)$
which implies $x\in A_{\tilde{g}}$ with $\left\lvert\nabla\tilde{g}\right\rvert^{+}(x)\leq 3\left\lvert\nabla g\right\rvert^{+}(x)$. Since $A_{g}\cap\overset{\circ}{C}=A_{\tilde{g}}\cap\overset{\circ}{C}$, where $\overset{\circ}{C}$ is the set of interior points of $C$, we have proved that $A_{g}\subseteq A_{\tilde{g}}$.
Now assume ${g}:{C}\rightarrow{M}$ is Lipschitz, then $A_{g}$ is a measurable subset of $C$ which is contained in $A_{\tilde{g}}$ where ${\tilde{g}}:{X}\rightarrow{M}$ is the extension found before. By assumption $\tilde{g}$ is a.e. $\mathcal{D}$-differentiable on $A_{\tilde{g}}$, then $g$ is a.e. $\mathcal{D}$-differentiable on $A_{g}$ and so $\mathcal{D}$ is complete. ∎